a particle initially at rest,start moving in a straight line with an acceleration a=6t+4m/S^.The distance covered by it in 3s is

Grade:11

1 Answers

Santhosh
6 years ago
On integrating once we get v=3t^2+4t+c. To find c put at t=o v=o because starts fron rest and again integrating s=t^3+2t^2 sub t=3we get 45mts